User Tools

Site Tools


operations:handover

Handover notes

Please use this page to write up any notes on anything that needs to be passed on to the next observer or should go into the end of experiment message.

Notes should be placed below the horizontal line as a new blog entry.

Continuing problems should be copied to the Current Issues page.

**Template for Handover notes:**

Be sure to add any specific notes about the experiment if required, for example “Special back end configuration for x reason”. Put as much detail as you can - the more the better.

It is ESSENTIAL to put your name after EVERY entry. Copy and paste the template into the new handover notes

====== Experiment Name ======

**Hobart 12m**

Disk VSN: abc-123 | Data volume at beginning: 2.34 GB

   * 1300UT a comment from me (user)

**Katherine 12m**

Disk VSN: | Data volume at beginning:  GB

   * 

**Yarragadee 12m**

Disk VSN: | Data volume at beginning:  GB

   * 1300UT Experiment started okay (JMc).

Format for entering a comment It is important to stick to this formula - there is a good reason we need this information, so please include it!

Time stamp in UT * Comment * (Your Name)

For Example:

20:00UT: Wind stows, Missed scans 012-3456 to 012-3457 (Jim)

20:50UT: Mark5 lost connection, had to restart Mark5 and reconfigure DBBC, called Jamie on call to fix, missed scans 012-3456 to 012-4567 (Jim)

Again - as much detail as you can. It is important to remember that these notes will be sent to the correlator, so please make sure they make sense to others.


AUM017

Hobart 12m

Disk VSN: flexbuffhb & HOB_0106

  • First good data from 189-1832. Config & mark5B issues prior to this time. (JMc)
  • 23:55 UT lost antenna comms (warren)
  • 01:49 UT antenna comms restored, probably not on source for two hours (w)

Katherine 12m

Disk VSN: flexbuffke

  • 1730UT Experiment started okay (Lim)

Yarragadee 12m

Disk VSN: flexbuffyg

  • 1730UT Experiment started okay (Lim)
2020/07/07 17:34 · Lim Chin Chuan
R1954

Ho

Disk VSN: HOB-1060/2000 | Data volume at beginning: 6604 gb.

  • Stuck drives?! Schedule resumed by scan 188-1823. Prior data invalid.
  • Received continuous error ma -4 device rx timed-out not response. At 10:16 UT terminate the PCFS and halt schedule. Received error “Allocating memory for antenna monitoring”. Resume schedule at 10:40 UT. Missed scans from 189-1015 through 189-1041 (mas)
  • about 1516 UT, 1603UT , Device rs timed-out on response (Lim)
2020/07/06 17:56 · Earl Lester
R4953

Yg

Disk VSN: USN-0181/2000 | Data volume at beginning: 0.353 GB

  • 1830UT Experiment started OK (Patrick)
  • ~0230 UT - Network issues, pings to pcfsyg varying up to ~10s causing ssh timeouts, etc. (JMc)
  • 1333UT mk5 connection closed. It might be a problem with the recorder or module, as restarting DIMino failed. The pointer was pointing to 0, fortunately, no data have been overridden per DirList before recover=2 issued. Missed scan 184-1348 to 184-1432. (Lim)
2020/07/01 17:43 · Patrick Yates
crf119

Calendar said this exp started at 4am, but it actually started at 3:30am. Yg started on time, But Hobart started ~15 mins late due to module change (and us thinking it did not start until 4 am)

Hobart26m

Disk VSN: HOB-1003 | Data volume at beginning: 0 GB

  • 1730UT Experiment started about ~15 mins late (Katie)
  • 2325 UT inserted checkmk5 procedure (Warren)
  • 1428UT FS crashed. Restarted FS & rebooted drive pc. Mk5 stuck in conflicting state, had to restart DIMino. Back on schedule & recording by 1520. Missed scans 183-1432 to 183-1516. (Patrick)

Yg

Disk VSN: HOB-1012 | Data volume at beginning: 0 GB

  • 1730UT Experiment started OK (Katie)
2020/06/30 17:38 · Katie Vandorou · 0 Comments
R1953

Hobart 26m

Disk VSN: HOB+0034 | Data volume at beginning: 3926 GB

  • 1700UT - Experiment started OK (Patrick)
  • 1038UT and 1138UT - ERROR ma -4 Device rx timed-out on response (Prad).
2020/06/29 16:39 · Patrick Yates
R4952

Disk VSN: USN-0017 | Data volume at beginning: 0 GB

Experiment Started OK. (Earl @ 4:30)

Logs properly set-up. (Earl @ 8:30)

2020/06/25 22:56 · Earl Lester · 0 Comments
/home/www/auscope/opswiki/data/pages/operations/handover.txt · Last modified: 2018/10/29 16:27 by Jesse Swan